This is actually a review word problem:
Two angles are complementary. One angle measures 17 degrees more than the other angle. Find the measure of the two angles.

Answer by nerdybill(7384)

Remember: When two angles are complementary, the sum of the two angles equal 90 degrees.
.
Let x = one of two angles
then from "One angle measures 17 degrees more than the other angle." we get
x+17 = second angle
.
x + x+17 = 90
2x + 17 = 90
2x = 73
x = 36.6 degrees
.
x+17 = 36.6+17 = 53.5 degrees
.
The two angles are
36.6 degrees and 53.5 degrees